OpenStudy (anonymous):

A lottery game asks you to pick six numbers from the range 1 through 40. How many ways are there to play this lottery game? Assume that this is a game where the order of your picks does NOT matter. 3,838,380 2,763,633,600 240 40

OpenStudy (anonymous):

since the order is not important we will use combination \[C_{6}^{40}=3,838,380\]

OpenStudy (ajspeller):

\[\left(\begin{matrix}40 \\ 6\end{matrix}\right)=\frac{ 40! }{ (40-6)!6! }\]

OpenStudy (anonymous):

ok ty